write an equation for the line perpendicular to y=2/3x+1 that contains (0,2)

Answer by richwmiller(17219)

parallel line would have the same slope
perpendicular line will have a slope that multiplied together is -1
so -3/2
y=-3/2x + b
2=-3/2(0)+b
2=b
y=-3/2x+2